. A lounge needs 72 m of carpet 4/5 m wide to cover it. How many meters ofcarpet,2/3m wide will cover the same lounge?
Answer by Alan3354(69443)

A lounge needs 72 m of carpet 4/5 m wide to cover it. How many meters of carpet, 2/3m wide will cover the same lounge?
============
With no dimensions given, only the area can be used.
---
72 * (4/5) = 72*0.8 = 57.6 sq meters
=============
L * (2/3) = 57.6
L = 3*57.6/2 = 86.4 meters